### v1

@@ 5.1 PLD and $f$-DP filters are NOT free in general
See Section˜A.1. ∎
A first implication Theorems˜2 and 3 is that there are no free $f$-DP filters in general:
Corollary 1 ($f$-DP filters are NOT free).
Consider $\mathcal{T}$ the family of all tradeoff curves, and $\mathcal{F}=\{f_{\varepsilon,\delta},\ \varepsilon\geq 0,\ \delta\in[0,1]\}$ the set of tradeoff curves for all $(\varepsilon,\delta)$-DP mechanisms.
Neither $\mathcal{T}$ nor $\mathcal{F}$ have free $f$-DP filters. Precisely, Algorithm˜2 is not $f$-DP.
@@
$\mathcal{F}\subset\mathcal{T}$, and $\mathcal{F}$ has crossing tradeoff functions, which are thus not well ordered. By Theorem˜3, Algorithm˜1 is not PLD for budget $B$, and so by Propositions˜10 and 11 Algorithm˜1 is not $f$-DP. ∎
Figure˜1 directly illustrates a counter-example to Theorem˜2 (Equation˜1) visually. We show that there exists tradeoff curves in $\mathcal{F}$ (approximate DP mechanisms) such that $\alpha\ \textrm{s.t.}\ g_{1}\otimes\operatorname{conv}(g_{1},g_{2})(\alpha)<\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})(\alpha)$. That is, the exists $f=g1$ and $\mathcal{H}=\{g_{1},g_{2}\}$ such that $f\otimes\operatorname{conv}(\mathcal{H})\neq\operatorname{conv}(f\otimes\mathcal{H})$, violating (Equation˜1). By Propositions˜10 and 11, there exists $L$ and $\mathcal{L}^{\prime}$ such that $L\oplus\sup{\mathcal{L}^{\prime}}\neq\sup\{L\oplus L^{\prime}:L^{\prime}\in\mathcal{L}^{\prime}\}$, violating Theorem˜2. In fact, Figure˜1 illustrates the two cases to account for, to directly prove a special case of Theorem˜3 for $(\varepsilon,\delta)$-DP:
Proposition 13 ($(\varepsilon,\delta)$-DP curves that cross always violate Theorem˜2 (Equation˜1)).
Consider $f\in\mathcal{T}$, and any set of tradeoff curves $\mathcal{F}$ such that $f_{\varepsilon_{1},\delta_{1}}$ and $f_{\varepsilon_{2},\delta_{2}}$ cross, and $f_{\varepsilon_{1},\delta_{1}},f_{\varepsilon_{2},\delta_{2}}\in\mathcal{F}$.
$\mathcal{F}$ does not have free $f$-DP filters. Precisely, Algorithm˜2 is not $f$-DP.
@@
Direct corollary of Theorem˜3, self standing direct proof in Section˜A.2. ∎
(a) Case 1: $g_{1}\otimes g_{1}$ and $g_{1}\otimes g_{2}$ cross.
(b) Case 2: $g_{1}\otimes g_{1}\geq g_{1}\otimes g_{2}$
Figure 1: Counter examples to Theorem 2 (Equation˜1), showing that $f$-DP filters are NOT free in general (for the set of all tradeoff curves $\mathcal{T}$), or when choosing adaptively among the set of $(\varepsilon,\delta)$-DP mechanisms. This is true even if $f\in\mathcal{F}$.
Another interesting case is that of pure DP, or $\varepsilon$-DP, mechanisms. Indeed, this family is well ordered ($\varepsilon_{1}>\varepsilon_{2}\geq 0\Rightarrow f_{\varepsilon_{1}}\preceq f_{\varepsilon_{2}}$, illustrated on Figure˜2(a)), and known to admit privacy filters under basic composition [RRUV16]. However, this family of DP mechanisms is not closed under $f$-DP or PLD composition, and its self composition is not well ordered anymore, as illustrated on Figure˜2(b). As we can see, this is sufficient to violate Theorems˜2 and 3, and hence rule out free $f$-DP/PLD filters for the family of $\varepsilon$-DP under $f$-DP/PLD composition.
(a) $\varepsilon$-DP admits two-step filters (example)
(b) $\varepsilon$-DP general filter counter-example
Figure 2: $\varepsilon$-DP mechanisms are well ordered, but not closed under composition. If $\mathcal{F}=\{f_{\varepsilon}:\ \varepsilon\geq 0\}$, then $\mathcal{F}^{2}\subset\mathcal{F}^{\infty}$ is not well ordered (Figure˜2(a)), and by Theorem˜3 pure DP mechanisms do not admit an $f$-DP/PLD filter (Figure˜2(b)).
We have shown that $f$-DP/PLD filters are not free in general (when $\mathcal{F}$ is the set of all tradeoff curves), or for the set of $(\varepsilon,\delta)$-DP mechanisms. These results are surprising, as all previous filter results were that filters are free (RDP/zCDP [feldman2021individual, lecuyer2021practical], GDP [koskelaindividual, smith2022fullyadaptivecompositiongaussian]) or asymptotically free ($(\varepsilon,\delta)$-DP [rogers2016privacy, whitehouse2023fully]). One might hope that restricting $f$, the bound on privacy loss used as input to the $f$-DP filter, would be sufficient to ensure free filters. Two natural candidates come to mind. First, restricting $f\in\mathcal{F}$, a practical condition of choosing privacy bounds in the set of tradeoff curves used in the filter. Second, we could restrict $f$ to a GDP curve $G_{\mu}$ (the tradeoff curve of a Gaussian mechanism), or an approximate GDP curve $f_{0,\delta}\otimes G_{\mu}$ to support approximate DP $f_{\varepsilon,\delta}$ tradeoff curves in $\mathcal{F}$. These are good candidates, as $\{G_{\mu}:\mu\in\mathbb{R}^{+}\}$ is the only known set of tradeoff curves to have free $f$-DP filters [koskelaindividual, smith2022fullyadaptivecompositiongaussian].
Unfortunately, neither restriction is sufficient:
Corollary 2.
Consider the set of $(\varepsilon,\delta)$-DP tradeoff curves $\mathcal{F}=\{f_{\varepsilon,\delta}:\varepsilon\geq 0,\ \delta\in[0,1]\}$. $\mathcal{F}$ does not have free $f$-DP filters even for $f\in\mathcal{F}$.

### v2

@@ 7.1.1 Useful Lemmata
Consider the Markov kernel $\kappa$ that takes in $(i,y)$ and outputs (i,1y)(i,1-y). It is obvious to see that $\kappa(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(0))=\kappa(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(1))$ and $\kappa(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(1))=\kappa(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(0))$. Thus, $(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(0),\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(1))$ dominates $(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(1),\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(0))$ and vice versa. In other words, $(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(0),\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(1))$ is symmetric.
The proof for $(\mathrm{RR}^{\downarrow}_{\mathbf{p},\bm{\varepsilon}}(0),\mathrm{RR}^{\downarrow}_{\mathbf{p},\bm{\varepsilon}}(1))$ is analogous. ∎
For the proofs below, it will be more convenient to deal with pairs rather than privacy loss. For this, we give an alternate definition of discretized privacy loss profile below.
Definition 14.
Let $(P,Q)$ be any pair of distributions. For any $0\leq\varepsilon<\varepsilon^{\prime}$, let
$I_{\varepsilon,\varepsilon^{\prime}}:=\begin{cases}[\varepsilon,\varepsilon^{\prime})&\text{ if }\varepsilon^{\prime}\neq\infty,\\ [\varepsilon,\varepsilon^{\prime})\cup\{\infty\}&\text{ if }\varepsilon^{\prime}=\infty.\end{cases}$
Moreover, let
$\Omega^{P,Q}_{\varepsilon,\varepsilon^{\prime}}=\left\{\omega~\middle|~\log\left(\frac{dP}{dQ}(\omega)\right)\in I_{\varepsilon,\varepsilon^{\prime}}\right\},$
and similarly
$\Omega^{Q,P}_{\varepsilon,\varepsilon^{\prime}}=\left\{\omega~\middle|~\log\left(\frac{dQ}{dP}(\omega)\right)\in I_{\varepsilon,\varepsilon^{\prime}}\right\}.$
The following observation follows directly from the definition of PLD.
Observation 13.
For any $L=\operatorname{PLD}(P,Q)$ and $\bm{\varepsilon}\in\mathcal{E}_{n}$, the discretized privacy loss profile of $L$ is equal to $\mathbf{p}=(p_{1},\dots,p_{n})$ where
$p_{i}=P{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}\cup\Omega^{Q,P}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}.$
The next lemma is a type of symmetry with regards to $\Omega^{P,Q}_{\varepsilon,\varepsilon^{\prime}}$ and $\Omega^{P,Q}_{\varepsilon^{\prime},\varepsilon}$:
Lemma 6.
For any symmetric pair $(P,Q)$ and $0\leq\varepsilon<\varepsilon^{\prime}$,
$P{\left({\Omega^{P,Q}_{\varepsilon,\varepsilon^{\prime}}}\right)}=Q{\left({\Omega^{Q,P}_{\varepsilon,\varepsilon^{\prime}}}\right)}$ and $P{\left({\Omega^{Q,P}_{\varepsilon,\varepsilon^{\prime}}}\right)}=Q{\left({\Omega^{P,Q}_{\varepsilon,\varepsilon^{\prime}}}\right)}.$
@@
We only prove the first equality as the second is analogous.
Let $F_{P,Q}$ and $F_{Q,P}$ denote the CDFs of $\operatorname{PLD}(P,Q)$ and $\operatorname{PLD}(Q,P)$, respectively. Since the hockey-stick curves $H_{x}(P\parallel Q)$ and $H_{x}(Q\parallel P)$ are equal, (Zhu et al., 2022, Lemma 23) implies that $F_{P,Q}$ and $F_{Q,P}$ are exactly the same. Thus, for $\varepsilon^{\prime}\neq\infty$ we have
$P{\left({\Omega^{P,Q}_{\varepsilon,\varepsilon^{\prime}}}\right)}=\lim_{\tilde{\varepsilon}\to(\varepsilon)^{-},\tilde{\varepsilon}^{\prime}\to(\varepsilon^{\prime})^{-}}F_{P,Q}(\tilde{\varepsilon}^{\prime})-F_{P,Q}(\tilde{\varepsilon})=\lim_{\tilde{\varepsilon}\to(\varepsilon)^{-},\tilde{\varepsilon}^{\prime}\to(\varepsilon^{\prime})^{-}}F_{Q,P}(\tilde{\varepsilon}^{\prime})-F_{Q,P}(\tilde{\varepsilon})=Q{\left({\Omega^{Q,P}_{\varepsilon,\varepsilon^{\prime}}}\right)}.$
For $\varepsilon^{\prime}=\infty$, we similarly have
$P{\left({\Omega^{P,Q}_{\varepsilon,\varepsilon^{\prime}}}\right)}=\lim_{\tilde{\varepsilon}\to(\varepsilon)^{-}}1-F_{P,Q}(\tilde{\varepsilon})=\lim_{\tilde{\varepsilon}\to(\varepsilon)^{-}}1-F_{Q,P}(\tilde{\varepsilon})=Q{\left({\Omega^{Q,P}_{\varepsilon,\varepsilon^{\prime}}}\right)}.$
7.1.2 Pessimistic Estimate Proof
We are now ready to prove Theorem˜11.
Proof of Theorem˜11.
Suppose that $L=\operatorname{PLD}(P,Q)$. Since both $(P,Q)$ and $(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(0),\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(1))$ are symmetric (Lemma˜5), by Lemma˜8, it suffices to show that $H_{x}(P\parallel Q)\leq H_{x}(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(0)\parallel\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(1))$ for any $x\geq 1$.
For $i\in[n]$, let $P_{i}$ (resp. $Q_{i}$) denote the distribution of $P$ (resp. $Q$) conditioned on the value being in $\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}\cup\Omega^{Q,P}_{\varepsilon_{i-1},\varepsilon_{i}}$. From Lemma˜6, we have that $P$ and $Q$ can be written as mixture distributions $\sum_{i\in[n]}p_{i}\cdot P_{i}$ and $\sum_{i\in[n]}p_{i}\cdot Q_{i}$, respectively. Since $P_{i}$’s support are disjoint and $Q_{i}$’s support are disjoint (and the supports of $P_{i},Q_{j}$ are disjoint for all $i\neq j$), we simply have
$H_{x}(P\parallel Q)=\sum_{i\in[n]}p_{i}\cdot H_{x}(P_{i}\parallel Q_{i}).$
Similarly, we have
$H_{x}(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(0)\parallel\mathrm{RR}_{\mathbf{p},\varepsilon}(1))=\sum_{i\in[n]}p_{i}\cdot H_{x}(\mathrm{RR}_{\varepsilon_{i}}(0)\parallel\mathrm{RR}_{\varepsilon_{i}}(1)).$
Thus, it suffices to show that [^3] This essentially follows from Kairouz et al. (2015) but we repeat the full proof here for completeness. $H_{x}(P_{i}\parallel Q_{i})\leq H_{x}(\mathrm{RR}_{\varepsilon_{i}}(0)\parallel\mathrm{RR}_{\varepsilon_{i}}(1))$ for all $i\in[n]$ and $x\geq 1$. To see this is true, consider three cases:
  • Case I: $x\geq e^{\varepsilon_{i}}$. Then, $H_{x}(P_{i}\parallel Q_{i})=0=H_{x}(\mathrm{RR}_{\varepsilon_{i}}(0)\parallel\mathrm{RR}_{\varepsilon_{i}}(1))$
  • Case II: $1\leq x\leq e^{\varepsilon_{i-1}}$. In this case, we have $H_{x}(\mathrm{RR}_{\varepsilon_{i}}(0)\parallel\mathrm{RR}_{\varepsilon_{i}}(1))=\frac{1-x\cdot e^{-\varepsilon_{i}}}{1+e^{-\varepsilon_{i}}}.$ Meanwhile, we have
    $H_{x}(P_{i}\parallel Q_{i})$ $=P_{i}{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}-x\cdot Q_{i}{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}$
    $=\frac{1}{p_{i}}\cdot\left(P{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}-x\cdot Q{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}\right)$
    $=\frac{P{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}-x\cdot Q{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}}{P{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}+P{\left({\Omega^{Q,P}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}}$
[…]
    $=\frac{x+1}{1+\frac{Q{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}}{P{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}}}-x$
    $\leq\frac{x+1}{1+e^{-\varepsilon_{i}}}-x$
    $=\frac{1-x\cdot e^{-\varepsilon_{i}}}{1+e^{-\varepsilon_{i}}},$
    where the inequality is due to the definition of $\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}$.
  • Case III: $x\in{\left({e^{\varepsilon_{i-1}},e^{\varepsilon_{i}}}\right)}$. Since $H_{x}$ is convex in $x$, we have
    $H_{x}(P_{i}\parallel Q_{i})$ $\leq\frac{x-e^{\varepsilon_{i-1}}}{e^{\varepsilon_{i}}-e^{\varepsilon_{i-1}}}\cdot H_{e^{\varepsilon_{i}}}(P_{i}\parallel Q_{i})+\frac{e^{\varepsilon_{i}}-x}{e^{\varepsilon_{i}}-e^{\varepsilon_{i-1}}}\cdot H_{e^{\varepsilon_{i-1}}}(P_{i}\parallel Q_{i})$
    $\leq\frac{x-e^{\varepsilon_{i-1}}}{e^{\varepsilon_{i}}-e^{\varepsilon_{i-1}}}\cdot H_{e^{\varepsilon_{i}}}(\mathrm{RR}_{\varepsilon_{i}}(0)\parallel\mathrm{RR}_{\varepsilon_{i}}(1))+\frac{e^{\varepsilon_{i}}-x}{e^{\varepsilon_{i}}-e^{\varepsilon_{i-1}}}\cdot H_{e^{\varepsilon_{i-1}}}(\mathrm{RR}_{\varepsilon_{i}}(0)\parallel\mathrm{RR}_{\varepsilon_{i}}(1))$
    $=H_{\alpha}(\mathrm{RR}_{\varepsilon_{i}}(0)\parallel\mathrm{RR}_{\varepsilon_{i}}(1)),$
    where the second inequality follows from Cases I and II above, and the last equality is due to linearity of $H_{x}(\mathrm{RR}_{\varepsilon_{i}}(0)\parallel\mathrm{RR}_{\varepsilon_{i}}(1))$ for $x\in I_{e^{\varepsilon_{i-1}},e^{\varepsilon_{i}}}$. ∎
7.1.3 Optimistic Estimate Proof
We next prove Theorem˜12. This proof is in fact simpler than the previous one and only two cases are required.
Proof of Theorem˜12.
Similar to the proof of Theorem˜11, it suffices to show that
$H_{x}(P_{i}\parallel Q_{i})\geq H_{x}(\mathrm{RR}_{\varepsilon_{i-1}}(0)\parallel\mathrm{RR}_{\varepsilon_{i-1}}(1))$
for all $i\in[n]$ and $x\geq 1$. To see this is true, consider two cases:
  • Case I: $x\geq e^{\varepsilon_{i-1}}$. In this case, we simply have $H_{x}(\mathrm{RR}_{\varepsilon_{i-1}}(0)\parallel\mathrm{RR}_{\varepsilon_{i-1}}(1))=0\leq H_{x}(P_{i}\parallel Q_{i})$.
  • Case II: $x\in[1,e^{\varepsilon_{i-1}})$. In this case, we have $H_{x}(\mathrm{RR}_{\varepsilon_{i-1}}(0)\parallel\mathrm{RR}_{\varepsilon_{i-1}}(1))=\frac{1-x\cdot e^{-\varepsilon_{i-1}}}{1+e^{-\varepsilon_{i-1}}}.$ Meanwhile, similar to Case II in the proof of Theorem˜11, we have
    $H_{x}(P_{i}\parallel Q_{i})=\frac{x+1}{1+\frac{Q{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}}{P{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}}}-x\geq\frac{x+1}{1+e^{-\varepsilon_{i-1}}}-x=\frac{1-x\cdot e^{-\varepsilon_{i-1}}}{1+e^{-\varepsilon_{i-1}}},$
    where the inequality is due to the definition of $\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}$. ∎
7.2 From PLD Filter to Additive Filter
Having established fundamental results on discretization, we can now state our “additive” privacy filter. In this filter, we have a budget $\bm{\ell}=(\ell_{1},\dots,\ell_{n})\in\mathbb{R}^{n}_{\geq 0}$ where informally $\ell_{i}$ corresponds to the budget for $\mathrm{RR}_{\varepsilon_{i}}$. When a mechanism comes, subtract its discretized privacy profile $\mathbf{p}$ from the budget. If the budget remains non-negative in all coordinates (denoted by [^4] Here we write $\mathbf{0}$ to denote the all-zero vector, and $\mathbf{u}\geq\mathbf{v}$ for vectors $\mathbf{u},\mathbf{v}$ if the inequality holds coordinate-wise. $\bm{\ell}\geq\mathbf{0}$), we subtract it from the budget and return the response to the adversary. This is formalized in Algorithm˜3.
Algorithm 3 Additive Discretized PL Profile Filter
0: Discretization points $\bm{\varepsilon}\in\mathcal{E}_{n}$, budgets $\bm{\ell}\in\mathbb{R}^{n}_{\geq 0}$, analyst $\mathcal{A}$, query capacity $k$, dataset $D$
for $t\leftarrow 1,\dots,k$ do
  $\mathcal{A}(Y_{1},\dots,Y_{t-1})$ gives mechanism $M_{t}$ that has PLD $L_{t}\in\mathcal{L}_{\mathrm{sym}}$
  $\mathbf{p}^{t}\leftarrow$ Discretized privacy loss profile of $L_{t}$ w.r.t. $\bm{\varepsilon}$
  $\bm{\ell}\leftarrow\bm{\ell}-\mathbf{p}^{t}$
  if $\bm{\ell}\geq\mathbf{0}$ then
   yield $Y_{t}\sim M_{t}(D)$
  else
   break
Setting of Parameters.
For the remainder of this section, we set the parameter $\bm{\varepsilon}$ as follows:
  • Let $n=1+\left\lceil\log_{2}{\left({\frac{k(\varepsilon+2)}{\varepsilon}}\right)}\right\rceil$.
  • Let $\varepsilon_{n}=\infty$
  • For all $i\in[n-1]$, let $\varepsilon_{i}=\frac{\varepsilon+2}{2^{n-1-i}}$.
Furthermore, let us define $\bm{\ell}^{*}=(\ell^{*}_{1},\dots,\ell^{*}_{n})$ as follows:
$\ell^{*}_{i}=\begin{cases}k&\text{ if }i=1\\ 8\cdot\max\left\{\ell\in\mathbb{N}~\middle|~H_{e^{\varepsilon}}(\mathrm{RR}_{\varepsilon_{i-1}}^{\otimes(\ell-1)}(0)\parallel\mathrm{RR}_{\varepsilon_{i-1}}^{\otimes(\ell-1)}(1))\leq 2\delta\right\}&\text{ if }1<i<n,\\ 4\delta&\text{ if }i=n\end{cases}$ (1)
Additive Profile Filter as Adversary.
Following notation in Sections˜3 and 5, we write the Additive Discretized PL Profile Filter (Algorithm˜3) can be viewed as the rule
$\operatorname{Discrete}_{\bm{\varepsilon},\bm{\ell}}^{k}(L_{1},\dots,L_{t})=\{L_{t+1}\in\mathcal{L}_{\mathrm{sym}}:\mathrm{disc}_{\bm{\varepsilon}}(L_{1})+\cdots+\mathrm{disc}_{\bm{\varepsilon}}(L_{t+1})\leq\bm{\ell}\}.$
Our main theorem in this subsection is that the Additive Discretized PL Profile Filter–when initialized with the above discretization points and privacy budgets–is less strict than the natural $(\varepsilon,\delta)$-DP PLD filter:
Theorem 14.
[…]
$\operatorname{Nat}_{\mathcal{L}_{\mathrm{sym}},R_{\varepsilon,\delta}}^{k}(L_{1},\dots,L_{t})\subseteq\operatorname{Discrete}_{\bm{\varepsilon},\bm{\ell}}^{k}(L_{1},\dots,L_{t}).$ (2)
@@ 7.2 From PLD Filter to Additive Filter
Consider any $L_{1},\dots,L_{t+1}\in\mathcal{L}_{\mathrm{sym}}$, and let $\mathbf{p}^{1}=\mathrm{disc}_{\bm{\varepsilon}}(L_{1}),\dots,\mathbf{p}^{t+1}=\mathrm{disc}_{\bm{\varepsilon}}(L_{t+1})$. Note that (2) is equivalent to showing the following: if $h_{L_{1}\oplus\cdots\oplus L_{t+1}}({e^{\varepsilon}})\leq\delta$, then we must have $p^{1}_{i}+\cdots+p^{t+1}_{i}\leq\ell^{*}_{i}$ for all $i\in[n]$. The case $i=1$ is obvious since $\ell^{*}_{1}=k\geq t$. We will henceforth only consider $i>1$.
We will show the contrapositive: Suppose that $p^{1}_{i}+\cdots+p^{t+1}_{i}>\ell^{*}_{i}$ for some $i\in\{2,\dots,n\}$. We will show that $h_{L_{1}\oplus\cdots\oplus L_{t+1}}({e^{\varepsilon}})>\delta$. Recall (from Theorem˜12) that $L_{j}$ dominates $(\mathrm{RR}^{\downarrow}_{\mathbf{p}^{j},\varepsilon}(0),\mathrm{RR}^{\downarrow}_{\mathbf{p}^{j},\varepsilon}(1))$. Thus, it suffices to show that
$H_{e^{\varepsilon}}\left(\mathrm{RR}^{\downarrow}_{\mathbf{p}^{1},\varepsilon}(0)\times\cdots\times\mathrm{RR}^{\downarrow}_{\mathbf{p}^{t+1},\varepsilon}(0)~\middle\|~\mathrm{RR}^{\downarrow}_{\mathbf{p}^{1},\varepsilon}(1)\times\cdots\times\mathrm{RR}^{\downarrow}_{\mathbf{p}^{t+1},\varepsilon}(1)\right)>\delta.$ (3)
For $X\in\{0,1\}$, consider post-processing $(y^{1},\dots,y^{t+1})\sim\mathrm{RR}^{\downarrow}_{\mathbf{p}^{1},\varepsilon}(X)\times\cdots\times\mathrm{RR}^{\downarrow}_{\mathbf{p}^{t+1},\varepsilon}(X)$ as follows, where $\tau\in\mathbb{N}$ will be specified below.
  • Recall that each $y^{j}=(y^{j}_{1},y^{j}_{2})$ is a tuple of two numbers: $y^{j}_{1}\in[n]$ and $y^{j}_{2}\in\{0,1\}$. Let $j_{1}<\dots<j_{m}$ be indices $j$ such that $y^{j}=i$.
  • If $m<\tau$, then output $\bot$. Otherwise, output $(y^{j_{1}},\dots,y^{j_{\tau}})$.
Let $R(X)$ denote the distribution of the output of the above process, and let $q$ denote the probability that the above process does not output $\bot$. (Note that $q$ does not depend on $X$.) Observe that, conditioned on not outputting $\bot$, the distribution $R(X)$ is exactly the same as $\mathrm{RR}_{\varepsilon_{i-1}}^{\otimes\tau}(X)$. Thus,
$H_{e^{\varepsilon}}\left(\mathrm{RR}^{\downarrow}_{\mathbf{p}^{1},\varepsilon}(0)\times\cdots\times\mathrm{RR}^{\downarrow}_{\mathbf{p}^{t+1},\varepsilon}(0)~\middle\|~\mathrm{RR}^{\downarrow}_{\mathbf{p}^{1},\varepsilon}(1)\times\cdots\times\mathrm{RR}^{\downarrow}_{\mathbf{p}^{t+1},\varepsilon}(1)\right)$
$\geq q\cdot H_{e^{\varepsilon}}\left(\mathrm{RR}_{\varepsilon_{i-1}}^{\otimes\tau}(0)~\middle\|~\mathrm{RR}_{\varepsilon_{i-1}}^{\otimes\tau}(1)\right).$ (4)
We will next consider two cases separately: $i=n$ and $1<i<n$.
Case I: $i=n$.
In this case, set $\tau=1$. Notice that $H_{e^{\varepsilon}}(\mathrm{RR}_{\varepsilon_{n-1}}^{\otimes\tau}(0)\parallel\mathrm{RR}_{\varepsilon_{n-1}}^{\otimes\tau}(1))=\frac{1-e^{\varepsilon-\varepsilon_{n-1}}}{1+e^{-\varepsilon_{n-1}}}$. Due to our choice $\varepsilon_{n-1}=\varepsilon+2$, we have $H_{e^{\varepsilon}}(\mathrm{RR}_{\varepsilon_{n-1}}^{\otimes\tau}(0)\parallel\mathrm{RR}_{\varepsilon_{n-1}}^{\otimes\tau}(1))>0.5$. Observe also that
$q$ $=1-(1-p^{1}_{n})\cdots(1-p^{t+1}_{n})\geq 1-e^{-p^{1}_{n}-\cdots-p^{t+1}_{n}}\geq 1-e^{-4\delta}\geq 2\delta,$
where the first and last inequalities follow from the well-known [^5] See e.g. Topsøe (2007). identity $\frac{x}{1+x}\leq\ln(1+x)\leq x$ for all x>1x>-1 (and from $\delta<1/4$). Combining the two inequalities and (4), we can conclude that (3) holds (for $i=n$) as desired.
Case II: $1<i<n$.
In this case, we set $\tau=\max\left\{\ell\in\mathbb{N}~\middle|~H_{e^{\varepsilon}}\left(\mathrm{RR}_{\varepsilon_{i-1}}^{\otimes(\ell-1)}(0)~\middle\|~\mathrm{RR}_{\varepsilon_{i-1}}^{\otimes(\ell-1)}(1)\right)\leq 2\delta\right\}$. By the choice of $\tau$, we have
$H_{e^{\varepsilon}}(\mathrm{RR}_{\varepsilon_{i-1}}^{\otimes\tau}(0)\parallel\mathrm{RR}_{\varepsilon_{i-1}}^{\otimes\tau}(1))>2\delta.$
We will next bound $q$. Observe that $m$ is distributed exactly as $Z=Z_{1}+\cdots+Z_{t+1}$ where $Z_{1}\sim\mathrm{Ber}(p^{1}_{i}),\dots,Z_{t+1}\sim\mathrm{Ber}(p^{t+1}_{i})$. Thus, we have $\mathop{\mathbb{E}}[Z]=p^{1}_{i}+\cdots+p^{t+1}_{i}>\ell^{*}_{i}$ and $\mathrm{Var}[Z]\leq\mathop{\mathbb{E}}[Z]$. Note that we set $\tau<\ell^{*}_{i}/2$. Thus, we can thus apply Chebyshev’s inequality to conclude that
$1-q=\Pr[Z<\tau]\leq\Pr[Z<\mathop{\mathbb{E}}[Z]/2]\leq\frac{\mathrm{Var}[Z]}{(\mathop{\mathbb{E}}[Z]/2)^{2}}\leq\frac{4}{\mathop{\mathbb{E}}[Z]}<\frac{4}{\ell^{*}_{i}}<\frac{1}{2}.$
Combining the two inequalities and (4), we can conclude that (3) holds as desired. ∎
7.3 Analyzing Additive Discretized PL Profile Filter
Due to Theorem˜14, it suffices to directly analyze the Additive Discretized PL Profile Filter. In particular, we will show the following:
Theorem 15.
For any $\varepsilon>0,\delta\in(0,1/4)$, let $\bm{\varepsilon},\bm{\ell}^{*}$ be as in (1). Then,
$\operatorname{PLD}{\left({\operatorname{Discrete}_{\bm{\varepsilon},\bm{\ell}^{*}}^{k}}\right)}\preceq R_{\chi\cdot(\varepsilon+1),\chi\cdot\delta},$
where
$\chi=O{\left({{\left({\log k+\log{\left({1+\frac{1}{\varepsilon}}\right)}}\right)}\cdot\log(1/\delta)}\right)}.$
Before we prove Theorem˜15, let us first note that it immediately implies the main theorem of this section (Theorem˜9).
Proof of Theorem˜9.
Combining Theorems˜14 and 15 respectively, we have
$\operatorname{PLD}{\left({\operatorname{Nat}_{\mathcal{L}_{\mathrm{sym}},R_{\varepsilon,\delta}}^{k}}\right)}\preceq\operatorname{PLD}{\left({\operatorname{Discrete}_{\bm{\varepsilon},\bm{\ell}^{*}}^{k}}\right)}\preceq R_{\chi\cdot(\varepsilon+1),\chi\cdot\delta}.$
[…]
$b^{*}_{i}=\begin{cases}0&\text{ if }i=n\\ 2\lceil\log(1/\delta)\rceil\cdot\ell^{*}_{i}&\text{ if }i\neq n.\end{cases}$ (5)
7.3.1 Useful Lemmata
We will also use the following version of “triangle inequality”. Note that the assumption here is stronger than just assumption that the TV distance between $P,P^{\prime}$ and $Q,Q^{\prime}$ are smaller than $\gamma$. This stronger assumption allows us to derive an additive bound that is only $\gamma$, rather than a weaker bound $(x+1)\cdot\gamma$ (if we were to only assume a bound on the TV distance). This is crucial for our proof of Theorem˜15, which requires a large value of $x$.
Lemma 7.
Let $x,\gamma\geq 0$, and let $P,Q$ be distributions on domain $\Omega$. Furthermore, let $P^{\prime},Q^{\prime}$ be distributions on domain $\Omega\cup\{\omega^{*}\}$ such that the following holds: $P(E)\geq P^{\prime}(E)$ and $Q(E)\geq Q^{\prime}(E)$ for all $E\subseteq\Omega$, and $P^{\prime}(\omega^{*})\leq\gamma$. Then, $H_{x}(P\parallel Q)\leq H_{x}(P^{\prime}\parallel Q^{\prime})+\gamma$.